OpenStudy (randomuser):

The art museum had a total of 224 visitors on Tuesday. Visitors older than 18 paid $12 for admission. Visitors 18 years or younger paid $8 for admission. The museum collected a total of $2,520. How many visitors were 18 or younger? 42 56 182 194

OpenStudy (randomuser):

I do not want the answer just the steps leading up to the answer

satellite73 (satellite73):

how many steps would you like? or rather, would you like to use two variables say \(x.y\) or just one ?

OpenStudy (randomuser):

we just guide me how to figure this problem out

satellite73 (satellite73):

it is not like there is one way given by god or your math teacher, there are many ways, that is why i asked did you want one variable or two?

OpenStudy (randomuser):

which ever on is easier

satellite73 (satellite73):

they are equivalent, so lets use two

satellite73 (satellite73):

say \(x\) is under 18, \(y\) is over 18, then what is \(x+y\) the total number of people who visited

OpenStudy (randomuser):

36

satellite73 (satellite73):

??

satellite73 (satellite73):

The art museum had a total of 224 visitors on Tuesday.

OpenStudy (randomuser):

oh i thought you asked my the sum of that equation?

satellite73 (satellite73):

lets try again x is the number of people under 18, y is the number older what is \(x+y\)?

OpenStudy (randomuser):

224

satellite73 (satellite73):

whew so one equation we can write down is \[x+y=224\] now we need another

satellite73 (satellite73):

if there are \(x\) people under 18 who pay $12 a person, how much do the under 18 people pay total (your answer needs and x in it)

OpenStudy (randomuser):

18x= the total of how much the 18 or younger payed for?

satellite73 (satellite73):

no, but close forget about the number 18, i think it is confusing you say they are red and green instead of over 18 and under 18

satellite73 (satellite73):

if there are red people who pay $12 each, then they pay 12x total

satellite73 (satellite73):

and if there are y green people who pay $8 each they pay 8y total

OpenStudy (randomuser):

I meant to put 8x = total

satellite73 (satellite73):

oh ok

satellite73 (satellite73):

ok typo there \[x+y=224\\ 8x+12y=2520\] thats better

OpenStudy (randomuser):

x = 182 y= 42 or is it opposite?

satellite73 (satellite73):

idk let me check

OpenStudy (randomuser):

i put it opposite

OpenStudy (randomuser):

x = 42 y = 182

satellite73 (satellite73):

yea x is 42, that is the answer they want, the younger

OpenStudy (randomuser):

the answer is 42

OpenStudy (randomuser):

I mean a

satellite73 (satellite73):

yes 42

OpenStudy (randomuser):

Thank you for your time and help

satellite73 (satellite73):

your welcome!